# The evolution and detection of vector superradiant instabilities

### v1

@@ Abstract
Ultralight vectors can extract energy and angular momentum from a Kerr black hole (BH) due to superradiant instability, resulting in the formation of a BH-condensate system. In this work, we carefully investigate the evolution of this system numerically with multiple superradiant modes. Simple formulas are obtained to estimate important timescales, maximum masses of different modes, as well as the BH mass and spin at various times. Due to the coexistence of modes with small frequency differences, the BH-condensate system emits gravitational waves with a unique beat signature, which could be directly observed by current and projected interferometers. Besides, the current BH spin-mass data from the binary BH merger events already excludes the vector mass in the range $5\times 10^{-15}\ \mathrm{eV}<\mu<9\times 10^{-12}\ \mathrm{eV}$.
@@ I Introduction
To detect the BH-condensate system, both direct and indirect methods are available. In terms of direct detection, the authors of Refs. Yoshino:2014wwa ; Isi:2018pzk ; Sun:2019mqb detail a targeted search for gravitational waves (GWs) using data from (Advanced) LIGO LIGOScientific:2014pky . To distinguish the system from other monochromatic GW sources, such as neutron stars, it has been suggested that the GWs from the system have a positive frequency drift due to self-gravity, while the GWs from neutron stars have a frequency drift in the opposite direction Arvanitaki:2014wva ; Baryakhtar:2017ngi . In a previous calculation with ultralight scalars Guo:2022mpr , we have shown that the subdominant modes with overtone $n>0$ could coexist for a long time with the dominant mode. With slightly different frequencies, the interference of these modes could produce a unique beat signature of the emitted GW. The beat period ranges from milliseconds to months, depending on the mass of the host BH. Explicit calculation also shows that the beat strength is within the capability of current and projected interferometers. Thus the beat signature is another fingerprint of the BH-condensate systems. In addition, unresolved GWs from BH-condensate systems can be studied as potential stochastic backgrounds Brito:2017wnc ; Brito:2017zvb .
@@
Ultralight vector fields, such as dark photons Holdom:1985ag ; Essig:2013lka , could also form such BH-condensate system via superradiance East:2013mfa ; Baryakhtar:2017ngi ; East:2017mrj ; East:2017ovw ; Cardoso:2018tly ; East:2018glu ; Isi:2018pzk ; Siemonsen:2019ebd . Existing studies mainly focus on the most unstable mode with overtone number $n=0$. The GW signal from such a BH-condensate system is then monochromatic with approximately constant amplitude within its lifetime. If the duration is long, it would be indistinguishable from other monochromatic GW sources, such as neutron stars. Similar to the scalar case, the subdominant modes could be important to break the degeneracy. In this work, we apply the analysis in Refs. Guo:2022mpr ; Cheng:2022jsw to ultralight vectors. Particularly, we focus on the parameter space where the Compton wavelengths of the vectors are comparable to the BH radius, i.e., $G_{\mathrm{N}}M\mu/(\hbar c)\sim\mathcal{O}(1)$, where $M$ is the mass of the BH and $\mu$ is the vector mass. This region of parameters has the strongest superradiant instability and is most relevant in phenomenology.
@@ II.1 The Proca equation in Kerr spacetime
The Kerr spacetime metric, characterized by mass $M$ and angular momentum $J$, has the following form in the Boyer-Lindquist coordinates Boyer:1966qh ,
@@
The inner horizon $r_{-}$ and outer horizon $r_{+}$ are,
@@
The Lagrangian of a free massive vector field $A_{\mu}$ is,
@@
where $\mu$ is the vector mass and,
$F_{\rho\sigma}\equiv\nabla_{\rho}A_{\sigma}-\nabla_{\sigma}A_{\rho},$ (7)
@@
is automatically satisfied. The corresponding stress-energy tensor of the vector field is,
@@
where we have ignored the imaginary part of the eigenfrequency $i\Gamma$ since it is much smaller than the real part $\omega$. This point will be evident below by comparing Eq. (20) to Eq. (23). With the Lorenz gauge condition, the time component of $A^{\rho}$ can be solved from its spatial components. The variables of the spatial components can be separated as,
@@
where the integers $n,\ j,\ l$ and $m$ represent the overtone number, the total angular momentum number, the azimuthal number, and the magnetic number, respectively. Another widely used quantity is the principal number $\bar{n}$, defined as $n+l+1$. In the rest of the paper, we will refer to a specific mode using a generalized spectroscopic notation ${}^{2s+1}L_{j,m}^{n}$, with $L=S,P,D,F\dots$ for $l=0,1,2,3$, respectively. When $l$ is undetermined, notation ${}^{2s+1}l_{j,m}^{n}$ is also used. In the presence of multiple modes, the vector field (11) can be expressed as,
@@
with $\tilde{r}\equiv 2r/(\bar{n}r_{\mathrm{B}})$ and $r_{\mathrm{B}}\equiv 1/(M\mu^{2})$ is analogous to the Bohr radius in the hydrogen atom. The angular piece $\mathbf{Y}_{l,jm}(\theta,\varphi)$ in Eq. (12) is the pure-orbital vector spherical harmonic function, which satisfies the eigenequation Thorne:1980ru ; Baryakhtar:2017ngi ,
@@
The above discussion assumes the non-relativistic limit. In the relativistic case, one could not simply separate the variables as in Eq. (12) Brito:2015oca . Previous efforts include a semi-analytical treatment under slow-rotation approximation Pani:2012bp ; Pani:2012vp ; Pani:2013pma and a numerical approach pioneered by Ref. Cardoso:2018tly , which obtains precise numerical results of the ${}^{3}S_{1,1}^{0}$ modes. In 2018, the authors of Ref. Frolov:2018ezx successfully separated the variables in the Proca equation with a general Kerr-NUT-(A)dS BH metric in all dimensions. Shortly after that, the obtained coupled equations in four-dimensional Kerr metric are solved both numerically and analytically Dolan:2018dqv ; Baumann:2019eav . Central to the variable-separating process is the ansatz,
@@
where $\mathcal{R}$ and $\mathcal{S}$ are the radial and angular functions, respectively. The polarization tensor $B^{\mu\nu}$ satisfies,
@@
where $\nu$ is a complex constant to be determined and $h_{\lambda\sigma}$ is the principal tensor Frolov:2017kze . More details of the calculation can be found in Refs. Frolov:2017kze ; Frolov:2018pys ; Frolov:2018ezx ; Dolan:2018dqv ; Baumann:2019eav .
@@ II.2 Superradiant rates
In this work, we employ the analytical results in the $M\mu\ll 1$ limit, first presented in Ref. Baumann:2019eav . The real part of the eigenfrequency with high-order correction is,
@@
where $\bar{n}=n+l+1$, $\alpha\equiv M\mu$ and,
@@
The imaginary part, which is also called the superradiance rate, has the following form,
@@
Figure 3: The superradiance rate $\Gamma$ as a function of the mass coupling $M\mu$ for the ${}^{3}S_{1,1}^{0}$ mode (upper panel) and the ${}^{3}P_{2,2}^{0}$ mode (lower panel). The curves of different colors represent different values of BH spin $a_{*}$. The solid curves are from the analytic approximation in Eq. (23), while the dotted curves are from numreical calculation using the method in Ref. Dolan:2018dqv .
Fig. 3 further shows the dependence of the superradiance rate of the ${}^{3}S_{1,1}^{0}$ and ${}^{3}P_{2,2}^{0}$ modes on BH spin $a_{*}$. With $a_{*}$ decreasing, the superradiance rate decreases, and the $M\mu$ range in which the rate is positive shrinks. For each value of $M\mu$, there is a critical BH spin below which the superradiance could not happen. In Fig. 3 we also compare the analytic approximation in Eq. (23) to the numerical results using the direct integration method in Ref. Dolan:2018dqv . It is expected that the error of Eq. (23) becomes larger at larger $M\mu$, since the analytical expression is effectively a Taylor expansion of $M\mu$. To have a sense of the difference, the relative error for the ${}^{3}S_{1,1}^{0}$ mode with $a_{*}=0.99$ is approximately $2.23\%$ at $M\mu=0.1$ and grows to roughly $90.0\%$ as $M\mu=0.4$.
@@ II.3 GW emission fluxes
In the case with only a single mode ${}^{3}l_{j,m}^{n}$, we follow the method described in Ref. Brito:2014wla and calculate the GW emission flux in the limit of $M\mu\ll 1$ using the Newman-Penrose (NP) formalism. An additional assumption of flat background metric is made to simplify the algebra. The GW emission flux of the ${}^{3}l_{j,m}^{n}$ mode at the leading order of $M\mu$ can be expressed as,
@@
For the dominant and subdominant modes with $j=1,2$, the corresponding coefficients are,
@@
The coefficient $C^{(0011)}_{\mathrm{GW}}$ has been calculated in Ref. Baryakhtar:2017ngi by solving the linear Einstein equation using the Green’s function. The estimated value is 6.4 in the flat background spacetime and 6060~{}6060 in the Schwarzschild geometry. By fitting the numerical results, the authors of Ref. Siemonsen:2019ebd obtain a value of $C^{(0011)}_{\mathrm{GW}}=16.66$. Our result is consistent with these previous calculations. Note that Eq. (26) is the radiation of each single mode, without the coannihilation of two vectors from different modes.
@@
Figure 4: The GW emission fluxes as functions of the mass coupling $M\mu$ for four superradiant modes ${}^{3}S_{1,1}^{0}$ (blue), ${}^{3}S_{1,1}^{1}$ (red), ${}^{3}P_{2,2}^{0}$ (green) and ${}^{3}P_{2,2}^{1}$ (orange). The solid curves are from Eq. (26) with Eqs. (27), assuming $M\mu\ll 1$ and no interference between different modes. The numerical results for the ${}^{3}S_{1,1}^{0}$ mode (blue dashed) Siemonsen:2019ebd and scalar ${}^{1}P_{1,1}^{0}$ mode (grey dashed) Guo:2022mpr are also shown for comparison.
@@
    Anniilaiton of a dominant vector and a subdominant vector, producing a graviton with frequency $\widetilde{\omega}_{3}=\omega_{d}+\omega_{s}$;
@@
These processes change the number of vectors in different modes, which are important in the time evolution of the BH-condensate. For $N_{d}\gg N_{s}$, the loss of the dominant-mode vectors is mainly via the first process. The loss of the subdominant-mode vectors is mainly via the third process. Using the stress tensor with two modes in the calculation explained in Ref. Brito:2014wla , one could obtain the GW flux from the third process. For the subdominant modes with $j=1$ and $j=2$, the results are,
@@
In the evolution of the BH-condensate below, we keep only the most important process for each mode. Specifically, we use Eqs. (26) for dominant modes and Eqs. (28) for subdominant modes.
The coexistence of multiple modes has another important consequence. Due to the small difference in the frequencies of the two modes, the dominant and the subdominant modes interfere, producing a unique beat signature in the GW emission flux. This beat signature from BH-scalar-condensate has been studied in detail in Ref. Guo:2022mpr . In this work, we study the similar signature of the BH-vector-condensate systems. We first replace the $\mathcal{T}_{\mu\nu}$ in Ref. Guo:2022mpr by the one with a Proca field,
@@
where the labels $i,j=d$ or $s$, representing the dominant mode or the subdominant mode respectively. The $\mathcal{A}^{\mu}_{(i)}$ is defined below Eq. (13) and $\mathcal{F}^{(\nu)}_{\mu\sigma}$ is defined by Eq. (7) with $A^{\mu}$ replaced by $\mathcal{A}^{\mu}_{(i)}$. Following the same steps in Ref. Guo:2022mpr , one obtains the emission flux when these two modes coexist,
@@
where $\widetilde{l}$ and $\widetilde{m}$ are the orbital and azimuthal numbers for the emitted GW. Specifically with the dominant mode (${}^{3}(j-1)_{j,j}^{0}$ and subdominant mode (${}^{3}(j-1)_{j,j}^{1}$ under consideration, one has $\widetilde{m}=2j$ and $\widetilde{l}\geq\widetilde{m}$. The phase angle $\widetilde{\phi}^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}$ is defined as $\arg\left(U^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}\right)$. Finally, the most important $U^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}$ are,
@@
$U_{22}^{(\widetilde{\omega}_{3})}\approx$ $\frac{4}{3}\sqrt{\frac{2\pi}{5}}(8+3i\pi)M^{4}\mu^{8},$ (33)
for $j=1$, and,
@@
$U_{44}^{(\widetilde{\omega}_{3})}\approx$ $-\frac{1}{135}\sqrt{\frac{\pi}{7}}(64+35i\pi)M^{6}\mu^{10}.$ (36)

### v2

# Evolution and detection of vector superradiant instabilities

@@ Abstract
Ultralight vectors can extract energy and angular momentum from a Kerr black hole (BH) due to superradiant instability, resulting in the formation of a BH-condensate system. In this work, we carefully investigate the evolution of this system numerically with multiple superradiant modes. Simple formulas are obtained to estimate important timescales, maximum masses of different modes, as well as the BH mass and spin at various times. Due to the coexistence of modes with small frequency differences, the BH-condensate system emits gravitational waves with a unique beat signature, which could be directly observed by current and projected interferometers. Besides, the current BH spin-mass data from the binary BH merger events already exclude the vector mass in the range $5\times 10^{-15}\ \mathrm{eV}<\mu<9\times 10^{-12}\ \mathrm{eV}$.
@@ I Introduction
To detect the BH-condensate system, both direct and indirect methods are available. In terms of direct detection, the authors of Refs. Yoshino:2014wwa ; Isi:2018pzk ; Sun:2019mqb detail a targeted search for GWs using data from the (Advanced) Laser Interferometer Gravitational Wave Observatory (LIGO) LIGOScientific:2014pky . To distinguish the system from other monochromatic GW sources, such as neutron stars, it has been suggested that the GWs from the system have a positive frequency drift due to self-gravity, while the GWs from neutron stars have a frequency drift in the opposite direction Arvanitaki:2014wva ; Baryakhtar:2017ngi . In a previous calculation with ultralight scalars Guo:2022mpr , we have shown that the subdominant modes with overtone $n>0$ could coexist for a long time with the dominant mode. With slightly different frequencies, the interference of these modes could produce a unique beat signature of the emitted GW. The beat period ranges from milliseconds to months, depending on the mass of the host BH. Explicit calculation also shows that the beat strength is within the capability of current and projected interferometers. Thus the beat signature is another fingerprint of the BH-condensate systems. In addition, unresolved GWs from BH-condensate systems can be studied as potential stochastic backgrounds Brito:2017wnc ; Brito:2017zvb .
@@
Ultralight vector fields, such as dark photons Holdom:1985ag ; Essig:2013lka , could also form such a BH-condensate system via superradiance East:2013mfa ; Baryakhtar:2017ngi ; East:2017mrj ; East:2017ovw ; Cardoso:2018tly ; East:2018glu ; Isi:2018pzk ; Siemonsen:2019ebd ; Chen:2022nbb ; Jia:2023see . Existing studies mainly focus on the most unstable mode with overtone number $n=0$. The GW signal from such a BH-condensate system is then monochromatic with approximately constant amplitude within its lifetime. If the duration is long, it would be indistinguishable from other monochromatic GW sources, such as neutron stars. Similar to the scalar case, the subdominant modes could be important to break the degeneracy. In this work, we apply the analysis in Refs. Guo:2022mpr ; Cheng:2022jsw to ultralight vectors. Particularly, we focus on the parameter space where the Compton wavelengths of the vectors are comparable to the BH radius, i.e., $G_{\mathrm{N}}M\mu/(\hbar c)\sim\mathcal{O}(1)$, where $M$ is the mass of the BH and $\mu$ is the vector mass. This region of parameters has the strongest superradiant instability and is most relevant in phenomenology.
@@ II.1 The Proca equation in Kerr spacetime
The Kerr spacetime metric, characterized by mass $M$ and angular momentum $J$, has the following form in the Boyer-Lindquist coordinates Boyer:1966qh :
@@
The inner horizon $r_{-}$ and outer horizon $r_{+}$ are
@@
The Lagrangian of a free massive vector field $A_{\mu}$ is
@@
where $\mu$ is the vector mass and
$F_{\rho\sigma}\equiv\nabla_{\rho}A_{\sigma}-\nabla_{\sigma}A_{\rho}$ (7)
@@
is automatically satisfied. The corresponding stress-energy tensor of the vector field is
@@
where we have ignored the imaginary part of the eigenfrequency $i\Gamma$ since it is much smaller than the real part $\omega$. This point will be evident below by comparing Eq. (20) to Eq. (23). With the Lorenz gauge condition, the time component of $A^{\rho}$ can be solved from its spatial components. The variables of the spatial components can be separated as
@@
where the integers $n,\ j,\ l$ and $m$ represent the overtone number, the total angular momentum number, the azimuthal number, and the magnetic number, respectively. Another widely used quantity is the principal number $\bar{n}$, defined as $n+l+1$. In the rest of the paper, we will refer to a specific mode using a generalized spectroscopic notation ${}^{2s+1}L_{j,m}^{n}$, with $L=S,P,D,F\dots$ for $l=0,1,2,3$, respectively. When $l$ is undetermined, the notation ${}^{2s+1}l_{j,m}^{n}$ is also used. In the presence of multiple modes, the vector field (11) can be expressed as
@@
with $\tilde{r}\equiv 2r/(\bar{n}r_{\mathrm{B}})$, and $r_{\mathrm{B}}\equiv 1/(M\mu^{2})$ is analogous to the Bohr radius in the hydrogen atom. The angular piece $\mathbf{Y}_{l,jm}(\theta,\varphi)$ in Eq. (12) is the pure-orbital vector spherical harmonic function, which satisfies the eigenequation Thorne:1980ru ; Baryakhtar:2017ngi ,
@@
The above discussion assumes the non-relativistic limit. In the relativistic case, one could not simply separate the variables as in Eq. (12) Brito:2015oca . Previous efforts include a semi-analytical treatment under slow-rotation approximation Pani:2012bp ; Pani:2012vp ; Pani:2013pma and a numerical approach pioneered by Ref. Cardoso:2018tly , which obtains precise numerical results of the ${}^{3}S_{1,1}^{0}$ modes. In 2018, inspired by the work of Ref. Lunin:2017drx , the authors of Ref. Frolov:2018ezx successfully separated the variables in the Proca equation with a general Kerr-NUT-(A)dS BH metric in all dimensions. Shortly after that, the obtained coupled equations in four-dimensional Kerr metric are solved both numerically and analytically Dolan:2018dqv ; Baumann:2019eav . Central to the variable-separating process is the ansatz
@@
where $\mathcal{R}$ and $\mathcal{S}$ are the radial and angular functions, respectively. The polarization tensor $B^{\mu\nu}$ satisfies
@@
where $\nu$ is a complex constant to be determined and $h_{\lambda\sigma}$ is the principal tensor Frolov:2017kze . More details of the calculation can be found in Refs. Lunin:2017drx ; Frolov:2017kze ; Frolov:2018pys ; Frolov:2018ezx ; Dolan:2018dqv ; Baumann:2019eav ; Lunin:2019pwz .
@@ II.2 Superradiant rates
In this work, we employ the analytical results in the $M\mu\ll 1$ limit, first presented in Ref. Baumann:2019eav . The real part of the eigenfrequency with high-order correction is
@@
where $\bar{n}=n+l+1$, $\alpha\equiv M\mu$ and
@@
The imaginary part, which is also called the superradiance rate, has the following form:
@@
Figure 3: The superradiance rate $\Gamma$ as a function of the mass coupling $M\mu$ for the ${}^{3}S_{1,1}^{0}$ mode (upper panel) and the ${}^{3}P_{2,2}^{0}$ mode (lower panel). The curves of different colors represent different values of BH spin $a_{*}$. The solid curves are from the analytic approximation in Eq. (23), while the dotted curves are from numerical calculation using the method in Ref. Dolan:2018dqv .
Figure 3 further shows the dependence of the superradiance rate of the ${}^{3}S_{1,1}^{0}$ and ${}^{3}P_{2,2}^{0}$ modes on BH spin $a_{*}$. With $a_{*}$ decreasing, the superradiance rate decreases, and the $M\mu$ range in which the rate is positive shrinks. For each value of $M\mu$, there is a critical BH spin below which the superradiance could not happen. In Fig. 3 we also compare the analytic approximation in Eq. (23) to the numerical results using the direct integration method in Ref. Dolan:2018dqv . It is expected that the error of Eq. (23) becomes larger at larger $M\mu$, since the analytical expression is effectively a Taylor expansion of $M\mu$. To have a sense of the difference, the relative error for the ${}^{3}S_{1,1}^{0}$ mode with $a_{*}=0.99$ is approximately $2.23\%$ at $M\mu=0.1$ and grows to roughly $90.0\%$ as $M\mu=0.4$.
@@ II.3 GW emission fluxes
In the case with only a single mode ${}^{3}l_{j,m}^{n}$, we follow the method described in Ref. Brito:2014wla and calculate the GW emission flux in the limit of $M\mu\ll 1$ using the Newman-Penrose (NP) formalism. An additional assumption of flat background metric is made to simplify the algebra. The GW emission flux of the ${}^{3}l_{j,m}^{n}$ mode at the leading order of $M\mu$ can be expressed as
@@
For the dominant and subdominant modes with $j=1,2$, the corresponding coefficients are
@@
The coefficient $C^{(0011)}_{\mathrm{GW}}$ has been calculated in Ref. Baryakhtar:2017ngi by solving the linear Einstein equation using the Green’s function. The estimated value is 6.4 in the flat background spacetime and $60$ in the Schwarzschild geometry. By fitting the numerical results, the authors of Ref. Siemonsen:2019ebd obtain a value of $C^{(0011)}_{\mathrm{GW}}=16.66$. Our result is consistent with these previous calculations. Note that Eq. (26) is the radiation of each single mode, without the coannihilation of two vectors from different modes.
@@
Figure 4: The GW emission fluxes as functions of the mass coupling $M\mu$ for four superradiant modes ${}^{3}S_{1,1}^{0}$ (blue), ${}^{3}S_{1,1}^{1}$ (red), ${}^{3}P_{2,2}^{0}$ (green) and ${}^{3}P_{2,2}^{1}$ (orange). The solid curves are from Eq. (26) with Eqs. (27), assuming $M\mu\ll 1$ and no interference between different modes. The numerical results for the ${}^{3}S_{1,1}^{0}$ mode (blue dashed) Siemonsen:2019ebd and scalar ${}^{1}P_{1,1}^{0}$ mode (gray dashed) Guo:2022mpr are also shown for comparison.
@@
    Annihilation of a dominant vector and a subdominant vector, producing a graviton with frequency $\widetilde{\omega}_{3}=\omega_{d}+\omega_{s}$;
@@
These processes change the number of vectors in different modes, which are important in the time evolution of the BH-condensate. For $N_{d}\gg N_{s}$, the loss of the dominant-mode vectors is mainly via the first process. The loss of the subdominant-mode vectors is mainly via the third process. Using the stress tensor with two modes in the calculation explained in Ref. Brito:2014wla , one could obtain the GW flux from the third process. For the subdominant modes with $j=1$ and $j=2$, the results are the following:
@@
In the evolution of the BH-condensate below, we keep only the most important process for each mode. Specifically, we use Eq. (26) for dominant modes and Eqs. (28) for subdominant modes.
The coexistence of multiple modes has another important consequence. Due to the small difference in the frequencies of the two modes, the dominant and the subdominant modes interfere, producing a unique beat signature in the GW emission flux. This beat signature from BH-scalar-condensate has been studied in detail in Ref. Guo:2022mpr . In this work, we study the similar signature of the BH-vector-condensate systems. We first replace $\mathcal{T}_{\mu\nu}$ in Ref. Guo:2022mpr by the one with a Proca field,
@@
where the labels $i,j=d$ or $s$, representing the dominant mode or the subdominant mode respectively. $\mathcal{A}^{\mu}_{(i)}$ is defined below Eq. (13), and $\mathcal{F}^{(\nu)}_{\mu\sigma}$ is defined by Eq. (7) with $A^{\mu}$ replaced by $\mathcal{A}^{\mu}_{(i)}$. Following the same steps in Ref. Guo:2022mpr , one obtains the emission flux when these two modes coexist,
@@
where $\widetilde{l}$ and $\widetilde{m}$ are the orbital and azimuthal numbers for the emitted GW, respectively. Specifically with the dominant mode (${}^{3}(j-1)_{j,j}^{0}$ and subdominant mode (${}^{3}(j-1)_{j,j}^{1}$ under consideration, one has $\widetilde{m}=2j$ and $\widetilde{l}\geq\widetilde{m}$. The phase angle $\widetilde{\phi}^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}$ is defined as $\arg\left(U^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}\right)$. Finally, the most important $U^{(\widetilde{\omega})}_{\widetilde{l}\widetilde{m}}$ are
@@
$U_{22}^{(\widetilde{\omega}_{3})}\approx$ $\frac{4}{3}\sqrt{\frac{2\pi}{5}}(8+3i\pi)M^{4}\mu^{8}$ (33)
for $j=1$, and
@@
$U_{44}^{(\widetilde{\omega}_{3})}\approx$ $-\frac{1}{135}\sqrt{\frac{\pi}{7}}(64+35i\pi)M^{6}\mu^{10}$ (36)